OpenStudy (anonymous):

How do I change y=3/5x-9 into standard form?

OpenStudy (anonymous):

write it in the form ax+by +c = 0 basic arithmetic manipulation.

OpenStudy (anonymous):

so I get -3/5x+y=-9, do I multiply both sides by negative nine to get a positive x?

OpenStudy (anonymous):

no. you add 9 to both sides.

jhonyy9 (jhonyy9):

3/5x -y -9=0 3-5xy-45x=0 45x +5xy -3 =0 5x(9+y) -3 =0